Jay is twelve years older than fay. If fay's age two years from now is half jay's age six years from now, how old will each be 10 years from now? Show algebraic solution
Answer by Theo(13342)

let j = jay's age now.
let f = fay's age now.

jay is currently 12 years older than fay.

you get j = f + 12

2 years from now, fay's age will be f + 2.
6 years from now, jay's age will be j + 6

you get f + 2 = 1/2 * (j + 6)

multiply both sides of this equation by 2 and you get 2f + 4 = j + 6

solve for j to get j = 2f - 2

you have j = f + 12 and j = 2f - 2

this means that f + 12 = 2f - 2 because both of them are equal to j.

solve for f to get f = 14

since j = f + 12, this means that j = 26.

jay is 26 years old now.
fay is 14 years old now.

in 2 years, fay will be 16.
in 6 years, jay will be 32.

16 = 1/2 * 31.
this means that fay'sage in 2 years will be half of jay's age in 6 years.

in 10 years, jay will be 36 and fay will be 24.

that should be your solution.